describe the principle of operation of thermosyphon hot water and explain how you would calculate the thermosyphon head and the efficiency of the system.
Answer #1
  • Thermosyphon is a passive heat exchange process, based on physical convection, that circulates a fluid without a mechanical pump being required. This process is being used in heating and cooling processes such as heat pumps, water heaters, boilers and furnaces for the diffusion of liquids and volatile gases. It also develops over gradients of air temperature such as those utilized in a wood fire or solar chimney.
  • The thermosyphon process describes that the cold water has a higher specific gravity than warm water and thus, being heavier is going to sink down. The collector is therefore often placed beyond the water storage tank, so that cold water from the tank enters the collector through a downstream water pipe.
  • 3-s2.0-B9780128037645000109-f10-01-97801Thermosiphon circulation reboiler.
  • Then the required circulation rate W of the thermosiphon cycle for the chosen vapor mass fraction x is determined.

    Calculation of the heat load Q for heating and evaporating: Q = Q(HEAT)+ Q(EVAP)= W*C*\DeltaHEAT+W+X*R

  • Required thermosiphon circulating rate W for the chosen evaporation rate x: W=Q/C* \Delta HEAT+X*R(kg/hr)